Conveyor belt # 1 can move 1,000 boxes in 8 minutes. Conveyor belt #2 can move 1,000 boxes in 7 minutes. If conveyor belt #3 is added and all three are used, the 1,000 boxes are moved in 3 minutes. How long would it take the third conveyor belt alone to do the same job? Round to the nearest whole number if necessary.
Answer by mananth(16946)


let us assume moving 1000 boxes is 1 job
Conveyor belt # 1 can move 1,000 boxes in 8 minutes.
so it does 1/8 of the job in 1 min
Conveyor belt #2 can move 1,000 boxes in 7 minutes.
In 1 minute it will move 1/7 of the job
If conveyor belt #3 is added and all three are used,
the 1,000 boxes are moved in 3 minutes.
in 1 minute all three move 1/3 of the job

let the third take x hours to do the job alone
it does 1/x of the job in 1 min
1/x+1/7+1/8 = 1/3
1/x = 1/3-(1/7+1/8)
1/x = 0.065
x = 15.27 hours